OpenStudy (anonymous):

n - (-45) = 90 i understand how to solve this but does the parentheses go first? because I could have 45 - -45 =90 just making sure thx

OpenStudy (anonymous):

n-45=90 n=2

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

its subtraction*

OpenStudy (asnaseer):

You are right - the parentheses go first, so you end up with: n - - 45 = 90 n + 45 = 90 n = 90 - 45 n = 45

OpenStudy (anonymous):

I am very smart, dont argue with me. The answer iis 2

OpenStudy (anonymous):

whos arguing ...and thx asnaseer

OpenStudy (asnaseer):

yw